# Statistically Valid Post-Deployment Monitoring Should Be Standard for AI-Based Digital Health

### v1

@@ 4.1 Covariate Shift
Following the definition in morenoTorresUnifyingViewDataset2012 , covariate shift corresponds to the case where $p_{t_{1}}(\mathbf{s},\mathbf{c})\neq p_{t_{0}}(\mathbf{s},\mathbf{c})$, while $\forall\left(\mathbf{s},\mathbf{c}\right)\in\mathcal{S}\times\mathcal{C},~{}p_{t_{1}}(y\mid\mathbf{s},\mathbf{c})=p_{t_{0}}(y\mid\mathbf{s},\mathbf{c})$. To formalize covariate shift detection as a statistical hypothesis test, let $\mathcal{D}^{sc}_{t_{0}}=\left\{\left(\mathbf{S},\mathbf{C}\right)_{i}\right\}_{i=1}^{n_{t_{0}}}$ and $\mathcal{D}^{sc}_{t_{1}}=\left\{\left(\mathbf{S},\mathbf{C}\right)_{i}\right\}_{i=1}^{n_{t_{1}}}$ denote two sets of i.i.d. data collected at $t_{0}$ and $t_{1}$. We denote $p_{t_{0}}\left(\mathbf{s},\mathbf{c}\right)$ and $p_{t_{1}}\left(\mathbf{s},\mathbf{c}\right)$ as the joint distributions of the input and clinical features at the pre- and post-deployment stages. The goal is to monitor whether a statistically significant change has occurred in this joint distribution of covariates $\left(\mathbf{S},\mathbf{C}\right)$, which can be posed as the following two-sample hypothesis testing problem:
@@
Under the null hypothesis $H_{0}$, the joint distribution of features and clinical variables remains unchanged between $t_{0}$ and $t_{1}$; that is, $p_{t_{0}}\left(\mathbf{s},\mathbf{c}\right)=p_{t_{1}}\left(\mathbf{s},\mathbf{c}\right),\forall\left(\mathbf{s},\mathbf{c}\right)\in\mathcal{S}\times\mathcal{C}$. The alternative hypothesis $H_{1}$ posits that $p_{t_{0}}\left(\mathbf{s},\mathbf{c}\right)\neq p_{t_{1}}\left(\mathbf{s},\mathbf{c}\right),\exists\left(\mathbf{s},\mathbf{c}\right)\in\mathcal{S}\times\mathcal{C}$, indicating a potential shift in the underlying distribution of patient covariates. We note that, for $H_{0}$ and $H_{1}$, and output label $y$, this test assumes that the conditional distribution of $y$ remains unchanged—that is, $p_{t_{0}}(y|\mathbf{s},\mathbf{c})=p_{t_{1}}(y|\mathbf{s},\mathbf{c})$ holds $\forall(\mathbf{s},\mathbf{c})\in\mathcal{S}\times\mathcal{C}$.
The samples $\mathcal{D}_{t_{0}}^{sc}$ and $\mathcal{D}_{t_{1}}^{sc}$ are compared using two-sample hypothesis tests to decide between $H_{0}$ and $H_{1}$. The choice of test depends on the monitoring objectives and assumptions about the data. A variety of parametric and non-parametric methods can be applied, which are reviewed in Appendix B.
@@ 4.2 Concept Drift
Following the definition in morenoTorresUnifyingViewDataset2012 , concept drift corresponds to: $p_{t_{1}}(y\mid\mathbf{s},\mathbf{c})\neq p_{t_{0}}(y\mid\mathbf{s},\mathbf{c}),\exists\left(\mathbf{s},\mathbf{c}\right)\in\mathcal{S}\times\mathcal{C}$ while $p_{t_{1}}(\mathbf{s},\mathbf{c})=p_{t_{0}}(\mathbf{s},\mathbf{c}),\forall\left(\mathbf{s},\mathbf{c}\right)\in\mathcal{S}\times\mathcal{C}$. We define $\mathcal{D}^{scy}_{t_{0}}=\{\left(\mathbf{s},\mathbf{c},y\right)_{i}\}_{i=1}^{n_{0}}$ and $\mathcal{D}^{scy}_{t_{1}}=\{\left(\mathbf{s},\mathbf{c},y\right)_{i}\}_{i=1}^{n_{1}}$ to represent the collection of covariates and true labels at $t_{0}$ and $t_{1}$. Furthermore, let $p_{t_{0}}\left(\mathbf{s},\mathbf{c},y\right)$ and $p_{t_{1}}\left(\mathbf{s},\mathbf{c},y\right)$ denote the joint distributions of $\left(\mathbf{S},\mathbf{C},Y\right)$ at $t_{0}$ and $t_{1}$, for which the pre- and post-deployment datasets $\mathcal{D}^{scy}_{t_{0}}$ and $\mathcal{D}^{scy}_{t_{1}}$ are sampled. We formulate the two-sample hypothesis testing problem to detect the concept drift as:
@@
Under $H_{0}$, the joint distribution of features, clinical/demographic and label variables remains consistent over time, indicating that the model’s performance has not degraded. Under $H_{1}$, the joint distribution of inputs and labels has changed, which may affect the model’s performance if it is sensitive to such distribution shifts. Similar to the covariate shift discussed in Section 4.1, the task of monitoring concept drift is framed as detecting distributional shifts. This involves comparing the datasets $\mathcal{D}_{t_{0}}^{scy}$ and $\mathcal{D}_{t_{1}}^{scy}$ using a two-sample test to determine whether to reject $H_{0}$. We refer readers to the two-sample tests described in Appendix B for further details.
OP: Choosing an Appropriate Test
A key open problem in concept drift detection is selecting an appropriate two-sample test for the joint distribution $p(\mathbf{s},\mathbf{c},y)$, particularly in the presence of mixed data types and limited labeled samples. While non-parametric methods like MMD borgwardt2006integrating , their performance can vary significantly depending on the dimensionality, feature types, and sample size—yet principled guidance for test selection remains lacking. Furthermore, in practice, the assumption that $p(\mathbf{s},\mathbf{c})$ remains fixed is often violated, making it difficult to attribute detected changes solely to concept drift rather than combined distributional shifts. Another key open challenge is handling label scarcity: since joint-distribution tests require labeled data, tests may suffer from low power in settings where only sparse or delayed feedback is available.
@@
Evaluating the hypothesis test requires post-deployment ground truth labels, however these are costly and time-consuming to obtain. To mitigate this, we need new approaches for surrogate models to approximate the true label $y$ in both pre- and post-deployment settings, building on the foundation of surrogate endpoints established in clinical trials prentice1989surrogate . In healthcare machine learning, surrogate labels are often derived from data correlated with clinical outcomes, such as billing codes hripcsak2014using , lab results yu2018predicting , or earlier outcomes like 30-day readmission used in place of 90-day outcomes liu2018reliable . Alternatively, surrogate labels can be generated from combinations of weak sources—heuristics, knowledge bases, or auxiliary models ratner2017snorkel ; wang2019weak . While practical in label-scarce settings, these proxies may introduce noise or degrade over time, limiting their reliability songLearningNoisyLabels2022 ; hagmannValidityProblemsClinical2023 ; zhouBriefIntroductionWeakly2018 . This can be formulated as a regression problem $y=J\left(\mathbf{s},\mathbf{c},\hat{y}\right)+\epsilon_{i}$, where $\hat{y}$ is model’s prediction, $y$ is a label, $\epsilon_{i}$ is the residual noise, and $J$ represents the function approximated by the surrogate model. $J$ can be estimated by the Prediction Aided by Surrogate Training (PAST) algorithm xiaPredictionAidedSurrogate2024 .
Beyond surrogate modeling, this setting motivates new directions in label-efficient hypothesis testing—an area that remains underexplored. For example, Li et al. li2022label propose a “bimodal" query strategy for active labeling of samples in two-sample tests. They show this preserves validity and enhances test power under label constraints.
@@ 5.1 Monitoring Performance Score
Direct assessment of the performance for an AI model $\hat{Y}=f(\mathbf{S},\mathbf{C}),\left(\mathbf{S},\mathbf{C}\right)\sim p\left(\mathbf{s},\mathbf{c}\right)$, is essential for detecting degradation. To do that, one needs to collect the true label $Y$ at the pre- and post-deployment stages, $t_{0}$ and $t_{1}$, and evaluate the performance score using them along with the prediction variable $\hat{Y}$. Specifically, we write $M=g\left(f,p\left(\mathbf{s},\mathbf{c},y\right)\right)$ to denote an evaluation function that outputs the performance score, or metric $M$, with respect to the model $f$ and the data distribution $p\left(\mathbf{s},\mathbf{c},y\right)$.
[…]
@@
The performance metrics and hypothesis tests presented above capture average performance over the entire distribution $p(\mathbf{s},\mathbf{c},y)$, overlooking significant performance variations across different subgroups. Inspired by Cohort Enrichment  [^2] Cohort Enrichment refers to identifying subsets of the data where a particular phenomenon (e.g., degradation) is amplified relative to the general population liu2020cohortenrichment and Exceptional Model Mining (EMM)  [^3] EMM is a generalization of subgroup discovery aimed at finding subpopulations where model behavior deviates significantly from the global norm—whether in performance, fairness, or other metrics leman2008emm strategies, one can systematically identify subgroups experiencing meaningful performance decline. We define the subgroup-specific performance as $M_{t}^{\mathcal{G}}=g\left(f,p_{t}\left(\mathbf{s},\mathbf{c},y\mid\mathcal{G}\right)\right)$ for a candidate subgroup $\mathcal{G}\subseteq\mathcal{S}\times\mathcal{C}$. The task of identifying subgroups with the most greatest performance degradation between $t_{0}$ and $t_{1}$ is then formalized as the following optimization problem: $\max_{\mathcal{G}\subseteq\mathcal{S}\times\mathcal{C}}M^{\mathcal{G}}_{t_{0}}-M^{\mathcal{G}}_{t_{1}}$, s.t. $\left|\mathcal{G}\right|\geq r$, where $r$ is a predefined minimum group size. Solving it offers valuable insights for: (1) identifying features where the model’s discriminative power has shifted, (2) detecting subgroups that experience disproportionate performance degradation, and (3) uncovering complex interaction patterns that may signal vulnerable populations.
@@
A separate potential direction on addressing label scarcity is rooted in active learning  cohn1996active ; gal2017deep ; hanneke2014theory ; balcan2006agnostic , which aims to develop classification models under limited label availability. In the context of our work, active learning offers a principled approach to selecting which covariate instances $(\mathbf{s},\mathbf{c})$ should be labeled, thereby improving the performance of the model $f$ efficiently. Typically, this involves constructing an acquisition function $q(\mathbf{s},\mathbf{c})$ that quantifies the informativeness of instances across the covariate space $\mathcal{S}\times\mathcal{C}$. The instance with the highest acquisition score, $\arg\max_{(\mathbf{s},\mathbf{c})\in\mathcal{S}\times\mathcal{C}}q(\mathbf{s},\mathbf{c})$, is selected for labeling and used to update the model $f$. Representative acquisition functions include ensemble-based uncertainty estimation techniques such as Query-by-Committee (QBC)seung1992query and deep ensembleslakshminarayanan2017simple . Intuitively, these functions measure the uncertainty of model predictions over the covariate space. Regions with the highest predictive uncertainty often correspond to areas where the model underperforms. Consequently, prioritizing label queries in these regions facilitates more effective detection of model degradation.
@@ 5.2 Monitoring Prediction Correctness
where $\hat{y}=f(\mathbf{s},\mathbf{c})$ denotes the model’s prediction. Herein, we also reuse the notations of pre- and post-deployment data, and define $\mathcal{D}^{scz}_{t_{0}}=\{\left(\mathbf{s},\mathbf{c},z\right)_{i}\}_{i=1}^{n_{0}}$ and $\mathcal{D}^{scz}_{t_{1}}=\{\left(\mathbf{s},\mathbf{c},z\right)_{i}\}_{i=1}^{n_{1}}$ to represent the collection of covariates and model correctness indicators at $t_{0}$ and $t_{1}$. Furthermore, let $p_{t_{0}}\left(\mathbf{s},\mathbf{c},z\right)$ and $p_{t_{1}}\left(\mathbf{s},\mathbf{c},z\right)$ denote the joint distributions of $\left(\mathbf{S},\mathbf{C},Z\right)$ at $t_{0}$ and $t_{1}$, for which the pre- and post-deployment datasets $\mathcal{D}^{scz}_{t_{0}}$ and $\mathcal{D}_{t_{1}}^{scz}$ are sampled. We formulate the following hypothesis test to detect distribution shifts in the model’s predictions:
@@
The nonparametric two-sample tests described in Appendix B for detecting distributional shifts can be applied to determine $H_{0}$ and $H_{1}$ by examining $\mathcal{D}_{t_{0}}^{scz}$ and $\mathcal{D}_{t_{1}}^{scz}$.
Post-deployment monitoring, framed as the task of detecting distributional shifts in prediction correctness, offers superior sensitivity to performance changes within local populations. This approach provides a distinct advantage over performance score monitoring, as discussed in Appendix 5.1. For instance, it can alert users to significant shifts in model performance within specific regions of the covariate space, even when the model’s overall performance remains stable.
@@
Similar to the open problems in the previous subsection, instead of focusing on the model’s overall performance, we can also identify subgroups responsible for performance decline in the case of joint distribution. By selecting a discrepancy function $\Delta(p_{t_{0}},p_{t_{1}})$—where $\Delta$ denotes a measure of discrepancy, such as an $f$-divergence renyi1961measures —to quantify the difference between distributions, we can formulate the task of identifying subgroups that exhibit distributional differences as the following optimization problem:$\max_{\mathcal{G}\subseteq\mathcal{S}\times\mathcal{C}}\Delta\left(p_{t_{0}}\left(\mathbf{s},\mathbf{c},z\mid\mathcal{G}\right),p_{t_{1}}\left(\mathbf{s},\mathbf{c},z\mid\mathcal{G}\right)\right)$, s.t. $\left|\mathcal{G}\right|\geq r$, where $r$ denotes the minimum size of the subgroup, pre-specified based on clinical considerations. Solving this optimization problem yields a subgroup $\mathcal{G}$ of size at least $r$ that maximizes the performance-related distributional discrepancy between $t_{0}$ and $t_{1}$.
@@
Lastly, performance drift often unfolds gradually rather than through abrupt shifts. Most existing hypothesis tests are calibrated for sharp deviations, making them less effective for detecting slow-onset concept drift. Designing statistically rigorous and sensitive tests that can detect such gradual degradation—especially within specific patient subgroups—remains an important open challenge. Promising directions include the use of active two-sample testing strategies berisha2024label , which adaptively select informative samples to boost test power under label scarcity, as well as adaptive windowing and monitoring methods davisDetectionCalibrationDrift2020 ; soin2022chexstray that track cumulative changes over time and are well-suited for detecting subtle shifts.

### v3

@@ 4.1 Covariate Shift
Following the definition in Moreno-Torres et al. (2012), covariate shift corresponds to the case where $p_{t_{1}}(\mathbf{s},\mathbf{c})\neq p_{t_{0}}(\mathbf{s},\mathbf{c})$, while $\forall\left(\mathbf{s},\mathbf{c}\right)\in\mathcal{S}\times\mathcal{C},~p_{t_{1}}(y\mid\mathbf{s},\mathbf{c})=p_{t_{0}}(y\mid\mathbf{s},\mathbf{c})$. To formalize covariate shift detection as a statistical hypothesis test, let $\mathcal{D}^{sc}_{t_{0}}=\left\{\left(\mathbf{S},\mathbf{C}\right)_{i}\right\}_{i=1}^{n_{t_{0}}}$ and $\mathcal{D}^{sc}_{t_{1}}=\left\{\left(\mathbf{S},\mathbf{C}\right)_{i}\right\}_{i=1}^{n_{t_{1}}}$ denote two sets of i.i.d. data collected at $t_{0}$ and $t_{1}$. We denote $p_{t_{0}}\left(\mathbf{s},\mathbf{c}\right)$ and $p_{t_{1}}\left(\mathbf{s},\mathbf{c}\right)$ as the joint distributions of the input and clinical features at the pre- and post-deployment stages. The goal is to monitor whether a statistically significant change has occurred in this joint distribution of covariates $\left(\mathbf{S},\mathbf{C}\right)$, which can be posed as the following two-sample hypothesis testing problem:
@@
We note that, for $H_{0}$ and $H_{1}$, and output label $y$, this test assumes that the conditional distribution of $y$ remains unchanged—that is, $p_{t_{0}}(y|\mathbf{s},\mathbf{c})=p_{t_{1}}(y|\mathbf{s},\mathbf{c})$ holds $\forall(\mathbf{s},\mathbf{c})\in\mathcal{S}\times\mathcal{C}$. Under the null hypothesis $H_{0}$, the joint distribution of features and clinical variables remains unchanged between $t_{0}$ and $t_{1}$; that is, $p_{t_{0}}\left(\mathbf{s},\mathbf{c}\right)=p_{t_{1}}\left(\mathbf{s},\mathbf{c}\right),\forall\left(\mathbf{s},\mathbf{c}\right)\in\mathcal{S}\times\mathcal{C}$. The alternative hypothesis $H_{1}$ posits that $p_{t_{0}}\left(\mathbf{s},\mathbf{c}\right)\neq p_{t_{1}}\left(\mathbf{s},\mathbf{c}\right),\exists\left(\mathbf{s},\mathbf{c}\right)\in\mathcal{S}\times\mathcal{C}$, indicating a potential shift in the underlying distribution of patient covariates.
The samples $\mathcal{D}_{t_{0}}^{sc}$ and $\mathcal{D}_{t_{1}}^{sc}$ are compared using two-sample hypothesis tests to decide between $H_{0}$ and $H_{1}$. The choice of test depends on the monitoring objectives and assumptions about the data. A variety of parametric and non-parametric methods can be applied, which are reviewed in Appendix D.
@@ 4.2 Concept Drift
Following the definition in Moreno-Torres et al. (2012), concept drift corresponds to: $p_{t_{1}}(y\mid\mathbf{s},\mathbf{c})\neq p_{t_{0}}(y\mid\mathbf{s},\mathbf{c}),\exists\left(\mathbf{s},\mathbf{c}\right)\in\mathcal{S}\times\mathcal{C}$ while $p_{t_{1}}(\mathbf{s},\mathbf{c})=p_{t_{0}}(\mathbf{s},\mathbf{c}),\forall\left(\mathbf{s},\mathbf{c}\right)\in\mathcal{S}\times\mathcal{C}$. We define $\mathcal{D}^{scy}_{t_{0}}=\{\left(\mathbf{S},\mathbf{C},Y\right)_{i}\}_{i=1}^{n_{0}}$ and $\mathcal{D}^{scy}_{t_{1}}=\{\left(\mathbf{S},\mathbf{C},Y\right)_{i}\}_{i=1}^{n_{1}}$ to represent the collection of covariates and true labels at $t_{0}$ and $t_{1}$. Furthermore, let $p_{t_{0}}\left(\mathbf{s},\mathbf{c},y\right)$ and $p_{t_{1}}\left(\mathbf{s},\mathbf{c},y\right)$ denote the joint distributions of $\left(\mathbf{S},\mathbf{C},Y\right)$ at $t_{0}$ and $t_{1}$, for which the pre- and post-deployment datasets $\mathcal{D}^{scy}_{t_{0}}$ and $\mathcal{D}^{scy}_{t_{1}}$ are sampled. We formulate the two-sample hypothesis testing problem to detect the concept drift as:
@@
We note that this test assumes that the joint distributions of the features and clinical/demographic variables at the pre- and post-deployment stages remain unchanged, that is $p_{t_{0}}(\mathbf{s},\mathbf{c})=p_{t_{1}}(\mathbf{s},\mathbf{c})$ holds $\forall(\mathbf{s},\mathbf{c})\in\mathcal{S}\times\mathcal{C}$. Under $H_{0}$, the joint distribution of features, clinical/demographic, and label variables remains consistent over time, indicating that the model’s performance has not degraded. Under $H_{1}$, the joint distribution of inputs and labels has changed, which may affect the model’s performance if it is sensitive to such distribution shifts. Similar to the covariate shift discussed in Section 4.1, the task of monitoring concept drift is framed as detecting distributional shifts. This involves comparing the datasets $\mathcal{D}_{t_{0}}^{scy}$ and $\mathcal{D}_{t_{1}}^{scy}$ using a two-sample test to determine whether to reject $H_{0}$. We refer readers to the two-sample tests described in Appendix D for further details.
OP: Drift Mechanism Attribution
While concept drift is formally defined as changes in $p(y\mid\mathbf{s},\mathbf{c})$, standard two-sample tests cannot directly test conditional distributions. We employ sequential testing Lipton et al. (2018): first test for label shift (Appendix B); if absent, test the joint distribution $p(\mathbf{s},\mathbf{c},y)$ for concept drift. When both shifts co-occur, distinguishing the mechanism requires additional methods. Class-conditional testing, separately test whether $p(\mathbf{s},\mathbf{c}\mid y)$ has changed for each outcome class, as label shift preserves these distributions while concept drift does not, though this requires sufficient labeled samples per class. Black Box Shift Estimation (BBSE) Lipton et al. (2018) can estimate the magnitude of label shift from unlabeled data, providing qualitative insight into the relative contribution of each mechanism. Alternative approaches like stratified binning or kernel methods Fukumizu et al. (2007) are possible but remain impractical due to the curse of dimensionality.
@@
Evaluating the hypothesis test requires post-deployment ground truth labels; however, these are costly and time-consuming to obtain. To mitigate this, we need new approaches for surrogate models to approximate the true label $y$ in both pre- and post-deployment settings, building on the foundation of surrogate endpoints established in clinical trials Prentice (1989). In healthcare machine learning, surrogate labels are often derived from data correlated with clinical outcomes, such as billing codes Hripcsak et al. (2014), lab results Yu et al. (2018), or earlier outcomes like 30-day readmission used in place of 90-day outcomes Liu et al. (2018). Alternatively, surrogate labels can be generated from combinations of weak sources—heuristics, knowledge bases, or auxiliary models Ratner et al. (2017); Wang et al. (2019). While practical in label-scarce settings, these proxies may introduce noise or degrade over time, limiting their reliability Song et al. (2022); Hagmann et al. (2023); Zhou (2018). This can be formulated as a regression problem $y=J\left(\mathbf{s},\mathbf{c},\hat{y}\right)+\epsilon_{i}$, where $\hat{y}$ is model’s prediction, $y$ is a label, $\epsilon_{i}$ is the residual noise, and $J$ represents the function approximated by the surrogate model. $J$ can be estimated by the Prediction Aided by Surrogate Training (PAST) algorithm Xia and Wainwright (2024). Beyond surrogate modeling, this setting motivates new directions in label-efficient hypothesis testing—an area that remains underexplored. For example, Li et al. Li et al. (2022, 2024b, 2024a) propose a query strategies for active labeling of samples in two-sample tests. They show this preserves validity and enhances test power under label constraints.
@@ 5.1 Monitoring Performance Score
Direct assessment of the performance for an AI model $\hat{Y}=f(\mathbf{S}),\left(\mathbf{S},\mathbf{C}\right)\sim p\left(\mathbf{s},\mathbf{c}\right)$, is essential for detecting degradation. To do that, one needs to collect the true label $Y$ at the pre- and post-deployment stages, $t_{0}$ and $t_{1}$, and evaluate the performance score using them along with the prediction variable $\hat{Y}$. Specifically, we write $M=g\left(f,p\left(\mathbf{s},\mathbf{c},y\right)\right)$ to denote an evaluation function that outputs the performance score, or metric $M$, with respect to the model $f$ and the data distribution $p\left(\mathbf{s},\mathbf{c},y\right)$.
There are many choices for the evaluation function $g$. Appendix C describes performance metrics that are commonly used. For instance, let $M_{t_{1}}$ (or $M_{t_{0}}$) denote the performance score of a model at $t_{1}$ (or $t_{0}$). If we select classification accuracy for $M_{t_{1}}=g\left(f,p_{t_{1}}\right)$, we then have $M_{t_{1}}=\int\int\int\mathbbm{1}_{f(\mathbf{s})=y}p_{t_{1}}\left(\mathbf{s},\mathbf{c},y\right)d\mathbf{s}d\mathbf{c}dy$. Typically, one does not have access to $p_{t_{0}}$ or $p_{t_{1}}$ to evaluate $M_{t_{0}}$ or $M_{t_{1}}$; instead, one resorts to computing the empirical metric, e.g., $\hat{M}_{t_{1}}=\frac{1}{\left|\mathcal{D}^{scy}_{t_{1}}\right|}\sum_{\left(\mathbf{s},\mathbf{c},y\right)\in\mathcal{D}^{scy}_{t_{1}}}\mathbbm{1}_{f(\mathbf{s})=y}$, where $\mathcal{D}^{scy}_{t_{1}}=\{\left(\mathbf{s},\mathbf{c},y\right)_{i}\}_{i=1}^{n_{1}}$.
@@
The performance metrics and hypothesis tests presented above capture average performance over the entire distribution $p(\mathbf{s},\mathbf{c},y)$, overlooking significant performance variations across different subgroups. Inspired by Cohort Enrichment  [^3] Cohort Enrichment refers to identifying subsets of the data where a particular phenomenon (e.g., degradation) is amplified relative to the general population Liu et al. (2020) and Exceptional Model Mining (EMM)  [^4] EMM is a generalization of subgroup discovery aimed at finding subpopulations where model behavior deviates significantly from the global norm—whether in performance, fairness, or other metrics Leman et al. (2008) strategies, one can systematically identify subgroups experiencing meaningful performance decline. We define the subgroup-specific performance as $M_{t}^{\mathcal{G}}=g\left(f,p_{t}\left(\mathbf{s},\mathbf{c},y\mid\mathcal{G}\right)\right)$ for a candidate subgroup $\mathcal{G}\subseteq\mathcal{S}\times\mathcal{C}$. The task of identifying subgroups with the most greatest performance degradation between $t_{0}$ and $t_{1}$ is then formalized as the following optimization problem: $\max_{\mathcal{G}\subseteq\mathcal{S}\times\mathcal{C}}M^{\mathcal{G}}_{t_{0}}-M^{\mathcal{G}}_{t_{1}}$, s.t. $\left|\mathcal{G}\right|\geq r$, where $r$ is a predefined minimum group size. Solving it offers valuable insights for: (1) identifying features where the model’s discriminative power has shifted, (2) detecting subgroups that experience disproportionate performance degradation, and (3) uncovering complex interaction patterns that may signal vulnerable populations.
@@
A separate potential direction on addressing label scarcity is rooted in active learning  Cohn et al. (1996); Gal et al. (2017); Hanneke and others (2014); Balcan et al. (2006), which aims to develop classification models under limited label availability. In the context of our work, active learning offers a principled approach to selecting which covariate instances $(\mathbf{s},\mathbf{c})$ should be labeled, thereby improving the performance of the model $f$ efficiently. Typically, this involves constructing an acquisition function $q(\mathbf{s},\mathbf{c})$ that quantifies the informativeness of instances across the covariate space $\mathcal{S}\times\mathcal{C}$. The instance with the highest acquisition score, $\arg\max_{(\mathbf{s},\mathbf{c})\in\mathcal{S}\times\mathcal{C}}q(\mathbf{s},\mathbf{c})$, is selected for labeling and used to update the model $f$. Representative acquisition functions include ensemble-based uncertainty estimation techniques such as Query-by-Committee (QBC)Seung et al. (1992) and deep ensemblesLakshminarayanan et al. (2017). Intuitively, these functions measure the uncertainty of model predictions over the covariate space. Regions with the highest predictive uncertainty often correspond to areas where the model underperforms. Consequently, prioritizing label queries in these regions facilitates more effective detection of model degradation.
@@ 5.2 Monitoring Prediction Correctness
where $\hat{y}=f(\mathbf{s})$ denotes the model’s prediction. Herein, we also reuse the notations of pre- and post-deployment data, and define $\mathcal{D}^{scz}_{t_{0}}=\{\left(\mathbf{S},\mathbf{C},Z\right)_{i}\}_{i=1}^{n_{0}}$ and $\mathcal{D}^{scz}_{t_{1}}=\{\left(\mathbf{S},\mathbf{C},Z\right)_{i}\}_{i=1}^{n_{1}}$ to represent the collection of covariates and model correctness indicators at $t_{0}$ and $t_{1}$. Furthermore, let $p_{t_{0}}\left(\mathbf{s},\mathbf{c},z\right)$ and $p_{t_{1}}\left(\mathbf{s},\mathbf{c},z\right)$ denote the joint distributions of $\left(\mathbf{S},\mathbf{C},Z\right)$ at $t_{0}$ and $t_{1}$, for which the pre- and post-deployment datasets $\mathcal{D}^{scz}_{t_{0}}$ and $\mathcal{D}_{t_{1}}^{scz}$ are sampled. We formulate the following hypothesis test to detect distribution shifts in the model’s predictions:
@@
The nonparametric two-sample tests described in Appendix G for detecting distributional shifts can be applied to evaluate the hypothesis from $\mathcal{D}_{t_{0}}^{scz}$ and $\mathcal{D}_{t_{1}}^{scz}$.
This approach provides a distinct advantage over average performance score monitoring, as discussed in Section 5.1. For instance, it can alert users to significant shifts in model performance within specific regions of the covariate space, even when the model’s overall performance remains stable.
@@
[…]
@@
Lastly, performance drift often unfolds gradually rather than through abrupt shifts. Given limited data, designing statistically rigorous and sensitive tests that can detect such gradual degradation—especially within specific patient subgroups—remains an important open challenge. Promising directions include the use of active two-sample testing strategies Li et al. (2024b), which adaptively select informative samples to boost test power under label scarcity, as well as adaptive windowing and monitoring methods Davis et al. (2020b); Soin et al. (2022) that track cumulative changes over time and are well-suited for detecting subtle shifts.